OpenStudy (anonymous):

if 3squae root 7/a square root of 7 equals 3 square root of 7/7 then a would equal?

OpenStudy (anonymous):

\[3\sqrt{7}/a \sqrt{7}equals 3\sqrt{7}/7 then a equals?\]